OpenStudy (anonymous):

What is the function of this table? x y 1 10 2 17 3 24 4 31 5 38 6 45

OpenStudy (anonymous):

y=7x+3

OpenStudy (anonymous):

omg thank you soo much! how do you find that out so fast? any tips you can give me please?

OpenStudy (anonymous):

each time x increases by one, y is increasing by 7. so the slope is 7, and it's a line. so it's y = 7x + b. Then use any of the datapoints to solve for b.

OpenStudy (anonymous):

thanks!! this webiste is awesome!

OpenStudy (anonymous):

slope = (y1-y2)/(x1-x2) at x=0 y=3 y = Sx + Y-intercept